y=|x+3|-2 I would really appreciate any help.
Answer by stanbon!(97)

Because of the "x" in the absolute this is a two-part
problem.
1st part:
If x>=-3, then y=(x+3)-2 = x+1
But, if x<-3, then y = -(x+3)-2 = -x-5.
The reason for this is |x| has two meanings as follows:
|x|=x if x>=0
|x|=-x if x<0
Cheers,
Stan H.